A wooden column has the form of a regular octagonal prism. Its base edge is 2. 25 meters and its height is 4. 75 meters. Find the weight of the column, Allowing a density of 0. 8 gram per cubic.
Answer by rothauserc(4718)

We want the volume of the column
:
Volume(V) = area of the regular octagon * its height
:
Area of a regular polygon = ((a^2)n) / (4*tan(360/2n)), where a is length of side, n is the number of sides
:
Area of regular octagon = ((2.25^2)6) / (4*tan(360/12)) = 13.1528
:
V = 13.1528 * 4.75 = 62.4758
:
************************************************
density = 0.8 gram per cubic what?
:
student answers - 0.8 gram per cm^3
:
our volume is 62.4758 m^3
:
there are 1 x 10^6 cm^3 in 1 m^3
:
therefore
:
62.4758 m^3 = 62,475,800 cm^3
:
weight = 62,475,800 * 0.8 = 49,980,640 grams
************************************************
: